AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 16 of 33  ELOG logo
ID Date Author Subject
  361   Wed Mar 2 03:40:40 2022 David, GeorgeTemperature Cycle

04:40
Target density is hovering between 2.0e-8 and 2.3e-8, sometimes dipping below 2.0e-8

We will conduct a temperature walkthrough while there is no beam,
04:40: setpoint 125K
04:46 setpoint 95K
04:56 setpoint 108K
04:59 setpoint 110K
05:00 setpoint 112K (here improved from 110K)
05:03 setpoint 114K
05:08 setpoint 113K

05:11
113K seems good, fluctuating between 2.3e-8 and 3.1e-8
will leave temperature here

  360   Wed Mar 2 01:24:48 2022 David, GeorgeIon Source Update

02:25
Zoran was unable to ignite the ion source.

He has given some parameters to the control room so they can wiggle stuff around and try and ignite it.
He also said it may at some point ignite itself.
In worse case scenario, the person who usually manages the ion source will come in from 7am-8am, read Zoran's notes and try and get it working again.

Also target density seems a little worse but still hanging above 2e-8 to 2e-4

  359   Tue Mar 1 23:43:15 2022 David, GeorgeBeam Intensity Drop

00:38
Intensity in the ring dropped by 2 orders of magnitude to 5e5 on injection

Lost rutherford rate (as seen in last e-log post)
All pressures seem fine

00:48
Called main control room, said they will take a look

00:59
Servo motor control restarted and homed

01:11
Ion source is confirmed to be off
Zoran will try and fix it, should take 1hr or so and not guaranteed to work.

 

  358   Tue Mar 1 23:30:04 2022 David, GeorgeCurrent status

00:30
Pressure in S1 slightly decreased from before but still high enough to measure data.
One new count at 20MeV

Target density at S1: 2e-8 - 2.7e-8
Pressures OK (one small spike just before 00:00 in YR08)
System Checks OK
White rabbit fail diff: 49
Data write speed: ~1.4Mb/s
Beam intensity at injection: ~7.7e7
Cryo Temp Setpoint: 110K

(Beam intensity in ring dropped as I was writing which is why there is no rutherford peak on exRate image)
 

Attachment 1: 2022-03-02_00-39-47xytotal.png
2022-03-02_00-39-47xytotal.png
Attachment 2: 2022-03-02_00-39-20exRate.png
2022-03-02_00-39-20exRate.png
Attachment 3: 2022-03-02_00-38-44extotal.png
2022-03-02_00-38-44extotal.png
  357   Tue Mar 1 22:10:26 2022 David, GeorgeCurrent status

Ring current 6.67e7 particles/s

pressure at S1: 2.3-2.8e-8

Merger, Tape server (WS3): 1574 kB/s, writing ok

Leakage current (WS4): 3.20 uA

DAQ (WS2): temperature checked, clock status: ok, adc calibration: ok, white rabbit check: failed

Online monitoring (WS 5): done

Attachment 1: 2022-03-01_23-14-43aida.png
2022-03-01_23-14-43aida.png
Attachment 2: 2022-03-01_23-13-57exrate.png
2022-03-01_23-13-57exrate.png
Attachment 3: 2022-03-01_23-13-40xyto.png
2022-03-01_23-13-40xyto.png
Attachment 4: 2022-03-01_23-13-10extot.png
2022-03-01_23-13-10extot.png
  356   Tue Mar 1 21:34:06 2022 PM, LVcheck

no change:

Check Clock Status': Passed 5, Failed 0

Check ADC Calibration: Passed 5, Failed 0

'Check the White rabbit decoder status: Passed 0, Failed 5

 

target density at S1: 2.2-2.7e-8

temp: 110K

beam at injection: 7e7

Attachment 1: Screenshot_from_2022-03-01_22-33-08.png
Screenshot_from_2022-03-01_22-33-08.png
Attachment 2: Screenshot_from_2022-03-01_22-32-53.png
Screenshot_from_2022-03-01_22-32-53.png
Attachment 3: Screenshot_from_2022-03-01_22-32-40.png
Screenshot_from_2022-03-01_22-32-40.png
Attachment 4: Screenshot_from_2022-03-01_22-32-30.png
Screenshot_from_2022-03-01_22-32-30.png
Attachment 5: Screenshot_from_2022-03-01_22-32-18.png
Screenshot_from_2022-03-01_22-32-18.png
Attachment 6: Screenshot_from_2022-03-01_22-32-04.png
Screenshot_from_2022-03-01_22-32-04.png
Attachment 7: Screenshot_from_2022-03-01_22-31-54.png
Screenshot_from_2022-03-01_22-31-54.png
Attachment 8: Screenshot_from_2022-03-01_22-31-34.png
Screenshot_from_2022-03-01_22-31-34.png
Attachment 9: Screenshot_from_2022-03-01_22-31-20.png
Screenshot_from_2022-03-01_22-31-20.png
Attachment 10: Screenshot_from_2022-03-01_22-31-10.png
Screenshot_from_2022-03-01_22-31-10.png
Attachment 11: Screenshot_from_2022-03-01_22-30-58.png
Screenshot_from_2022-03-01_22-30-58.png
Attachment 12: Screenshot_from_2022-03-01_22-30-46.png
Screenshot_from_2022-03-01_22-30-46.png
Attachment 13: Screenshot_from_2022-03-01_22-30-31.png
Screenshot_from_2022-03-01_22-30-31.png
Attachment 14: Screenshot_from_2022-03-01_22-30-09.png
Screenshot_from_2022-03-01_22-30-09.png
  355   Tue Mar 1 20:41:28 2022 PM, LVcheck

no change:

Check Clock Status': Passed 5, Failed 0

Check ADC Calibration: Passed 5, Failed 0

'Check the White rabbit decoder status: Passed 0, Failed 5

 

target density at S1: 2.3-2.7e-8

temp: 110K

beam at injection: 7e7

Attachment 1: Screenshot_from_2022-03-01_21-41-02.png
Screenshot_from_2022-03-01_21-41-02.png
Attachment 2: Screenshot_from_2022-03-01_21-40-45.png
Screenshot_from_2022-03-01_21-40-45.png
Attachment 3: Screenshot_from_2022-03-01_21-40-33.png
Screenshot_from_2022-03-01_21-40-33.png
Attachment 4: Screenshot_from_2022-03-01_21-40-19.png
Screenshot_from_2022-03-01_21-40-19.png
Attachment 5: Screenshot_from_2022-03-01_21-40-03.png
Screenshot_from_2022-03-01_21-40-03.png
Attachment 6: Screenshot_from_2022-03-01_21-39-35.png
Screenshot_from_2022-03-01_21-39-35.png
Attachment 7: Screenshot_from_2022-03-01_21-39-20.png
Screenshot_from_2022-03-01_21-39-20.png
Attachment 8: Screenshot_from_2022-03-01_21-39-00.png
Screenshot_from_2022-03-01_21-39-00.png
Attachment 9: Screenshot_from_2022-03-01_21-38-50.png
Screenshot_from_2022-03-01_21-38-50.png
Attachment 10: Screenshot_from_2022-03-01_21-38-31.png
Screenshot_from_2022-03-01_21-38-31.png
Attachment 11: Screenshot_from_2022-03-01_21-38-06.png
Screenshot_from_2022-03-01_21-38-06.png
Attachment 12: Screenshot_from_2022-03-01_21-37-50.png
Screenshot_from_2022-03-01_21-37-50.png
Attachment 13: Screenshot_from_2022-03-01_21-37-40.png
Screenshot_from_2022-03-01_21-37-40.png
Attachment 14: Screenshot_from_2022-03-01_21-37-27.png
Screenshot_from_2022-03-01_21-37-27.png
  354   Tue Mar 1 19:41:24 2022 PM, LVtarget temperature walkthrough

A target temperature walkthrough has been initiated at 20:41 since the target density at S1 had slightly decreased (with ~0.1-0.2)

Current temprature 116K

set to 126K at 20:41

set to 95K at 20:47:30

set to 108.5K at 21:01:30, at this temperature was a sweetspot observerd for target density.

set to 109.7 at 21:08:30

set to 110K at 21:13:30 - we keep at this value for a while. the values at S1 are changing between 2.2-2.9e-8

  353   Tue Mar 1 19:40:28 2022 PM, LVcheck

no change:

Check Clock Status': Passed 5, Failed 0

Check ADC Calibration: Passed 5, Failed 0

'Check the White rabbit decoder status: Passed 0, Failed 5

 

target density at S1: 2.1-2.3e-8

temp: 116K

beam at injection: 7e7

Attachment 1: Screenshot_from_2022-03-01_20-39-20.png
Screenshot_from_2022-03-01_20-39-20.png
Attachment 2: Screenshot_from_2022-03-01_20-39-06.png
Screenshot_from_2022-03-01_20-39-06.png
Attachment 3: Screenshot_from_2022-03-01_20-38-50.png
Screenshot_from_2022-03-01_20-38-50.png
Attachment 4: Screenshot_from_2022-03-01_20-38-35.png
Screenshot_from_2022-03-01_20-38-35.png
Attachment 5: Screenshot_from_2022-03-01_20-38-22.png
Screenshot_from_2022-03-01_20-38-22.png
Attachment 6: Screenshot_from_2022-03-01_20-37-52.png
Screenshot_from_2022-03-01_20-37-52.png
Attachment 7: Screenshot_from_2022-03-01_20-37-42.png
Screenshot_from_2022-03-01_20-37-42.png
Attachment 8: Screenshot_from_2022-03-01_20-37-30.png
Screenshot_from_2022-03-01_20-37-30.png
Attachment 9: Screenshot_from_2022-03-01_20-37-16.png
Screenshot_from_2022-03-01_20-37-16.png
Attachment 10: Screenshot_from_2022-03-01_20-37-00.png
Screenshot_from_2022-03-01_20-37-00.png
Attachment 11: Screenshot_from_2022-03-01_20-36-41.png
Screenshot_from_2022-03-01_20-36-41.png
Attachment 12: Screenshot_from_2022-03-01_20-36-28.png
Screenshot_from_2022-03-01_20-36-28.png
  352   Tue Mar 1 18:43:41 2022 PM, LVcheck

no change:

Check Clock Status': Passed 5, Failed 0

Check ADC Calibration: Passed 5, Failed 0

'Check the White rabbit decoder status: Passed 0, Failed 5

 

target density at S1: 2.1-2.4e-8

temp: 116K

Attachment 1: Screenshot_from_2022-03-01_19-42-59.png
Screenshot_from_2022-03-01_19-42-59.png
Attachment 2: Screenshot_from_2022-03-01_19-42-45.png
Screenshot_from_2022-03-01_19-42-45.png
Attachment 3: Screenshot_from_2022-03-01_19-41-54.png
Screenshot_from_2022-03-01_19-41-54.png
Attachment 4: Screenshot_from_2022-03-01_19-41-39.png
Screenshot_from_2022-03-01_19-41-39.png
Attachment 5: Screenshot_from_2022-03-01_19-41-25.png
Screenshot_from_2022-03-01_19-41-25.png
Attachment 6: Screenshot_from_2022-03-01_19-41-12.png
Screenshot_from_2022-03-01_19-41-12.png
Attachment 7: Screenshot_from_2022-03-01_19-41-00.png
Screenshot_from_2022-03-01_19-41-00.png
Attachment 8: Screenshot_from_2022-03-01_19-40-29.png
Screenshot_from_2022-03-01_19-40-29.png
Attachment 9: Screenshot_from_2022-03-01_19-40-19.png
Screenshot_from_2022-03-01_19-40-19.png
Attachment 10: Screenshot_from_2022-03-01_19-40-00.png
Screenshot_from_2022-03-01_19-40-00.png
Attachment 11: Screenshot_from_2022-03-01_19-39-40.png
Screenshot_from_2022-03-01_19-39-40.png
Attachment 12: Screenshot_from_2022-03-01_19-39-26.png
Screenshot_from_2022-03-01_19-39-26.png
Attachment 13: Screenshot_from_2022-03-01_19-39-08.png
Screenshot_from_2022-03-01_19-39-08.png
Attachment 14: Screenshot_from_2022-03-01_19-38-49.png
Screenshot_from_2022-03-01_19-38-49.png
  351   Tue Mar 1 17:36:04 2022 PM, LVcheck

no change:

Check Clock Status': Passed 5, Failed 0

Check ADC Calibration: Passed 5, Failed 0

'Check the White rabbit decoder status: Passed 0, Failed 5

 

target density at S1: 2.1-2.4e-8

temp: 116K

Attachment 1: Screenshot_from_2022-03-01_18-43-15.png
Screenshot_from_2022-03-01_18-43-15.png
Attachment 2: Screenshot_from_2022-03-01_18-42-59.png
Screenshot_from_2022-03-01_18-42-59.png
Attachment 3: Screenshot_from_2022-03-01_18-42-21.png
Screenshot_from_2022-03-01_18-42-21.png
Attachment 4: Screenshot_from_2022-03-01_18-41-49.png
Screenshot_from_2022-03-01_18-41-49.png
Attachment 5: Screenshot_from_2022-03-01_18-41-30.png
Screenshot_from_2022-03-01_18-41-30.png
Attachment 6: Screenshot_from_2022-03-01_18-41-04.png
Screenshot_from_2022-03-01_18-41-04.png
Attachment 7: Screenshot_from_2022-03-01_18-40-39.png
Screenshot_from_2022-03-01_18-40-39.png
Attachment 8: Screenshot_from_2022-03-01_18-40-07.png
Screenshot_from_2022-03-01_18-40-07.png
Attachment 9: Screenshot_from_2022-03-01_18-39-34.png
Screenshot_from_2022-03-01_18-39-34.png
Attachment 10: Screenshot_from_2022-03-01_18-39-13.png
Screenshot_from_2022-03-01_18-39-13.png
Attachment 11: Screenshot_from_2022-03-01_18-38-57.png
Screenshot_from_2022-03-01_18-38-57.png
Attachment 12: Screenshot_from_2022-03-01_18-38-41.png
Screenshot_from_2022-03-01_18-38-41.png
Attachment 13: Screenshot_from_2022-03-01_18-35-09.png
Screenshot_from_2022-03-01_18-35-09.png
Attachment 14: Screenshot_from_2022-03-01_18-34-35.png
Screenshot_from_2022-03-01_18-34-35.png
  350   Tue Mar 1 16:57:34 2022 CB, LV, PMHNew run

Beam issues due to too long storage time in the last hours. Motors moved out and not moving.

Decreased storage time to 70 s. Seems OK now. Restarted motor movement as before

Started new run R86

  349   Tue Mar 1 14:01:01 2022 MS, SHBcheck

'Check Clock Status', 'Check ADC Calibration' and 'Check the White rabbit decoder status' : no change

Attachment 1: Screenshot_from_2022-03-01_15-00-30.png
Screenshot_from_2022-03-01_15-00-30.png
Attachment 2: Screenshot_from_2022-03-01_15-00-17.png
Screenshot_from_2022-03-01_15-00-17.png
Attachment 3: Screenshot_from_2022-03-01_14-58-01.png
Screenshot_from_2022-03-01_14-58-01.png
Attachment 4: Screenshot_from_2022-03-01_14-57-50.png
Screenshot_from_2022-03-01_14-57-50.png
Attachment 5: Screenshot_from_2022-03-01_14-57-22.png
Screenshot_from_2022-03-01_14-57-22.png
Attachment 6: Screenshot_from_2022-03-01_14-56-49.png
Screenshot_from_2022-03-01_14-56-49.png
Attachment 7: Screenshot_from_2022-03-01_14-56-35.png
Screenshot_from_2022-03-01_14-56-35.png
Attachment 8: Screenshot_from_2022-03-01_14-56-24.png
Screenshot_from_2022-03-01_14-56-24.png
Attachment 9: Screenshot_from_2022-03-01_14-56-10.png
Screenshot_from_2022-03-01_14-56-10.png
Attachment 10: Screenshot_from_2022-03-01_14-55-49.png
Screenshot_from_2022-03-01_14-55-49.png
Attachment 11: Screenshot_from_2022-03-01_14-55-40.png
Screenshot_from_2022-03-01_14-55-40.png
  348   Tue Mar 1 12:58:32 2022 CB, MS, SS, MLSynced pneumatic motion

New run R84

Attempting to sync pneumatic movement with ring. Still see something on t out -> p out

Command to move out was being given relative to the wrong time.

Fixed it. Still see one-two events every time we move out around 4-6 MeV. Attempted to greatly increase time before next injection - no difference. Microphonics from moving the detectors? It's just one-two counts on entire detector, acceptable.

 

New run R85

Now moving in/out

Potentiometer: 23.10 mm to 67.19 mm

  347   Tue Mar 1 12:33:21 2022 MS, SHBcheck

'Check Clock Status', 'Check ADC Calibration' and 'Check the White rabbit decoder status': no change

Attachment 1: Screenshot_from_2022-03-01_13-40-22.png
Screenshot_from_2022-03-01_13-40-22.png
Attachment 2: Screenshot_from_2022-03-01_13-40-07.png
Screenshot_from_2022-03-01_13-40-07.png
Attachment 3: Screenshot_from_2022-03-01_13-39-50.png
Screenshot_from_2022-03-01_13-39-50.png
Attachment 4: Screenshot_from_2022-03-01_13-39-33.png
Screenshot_from_2022-03-01_13-39-33.png
Attachment 5: Screenshot_from_2022-03-01_13-39-03.png
Screenshot_from_2022-03-01_13-39-03.png
Attachment 6: Screenshot_from_2022-03-01_13-38-05.png
Screenshot_from_2022-03-01_13-38-05.png
Attachment 7: Screenshot_from_2022-03-01_13-37-47.png
Screenshot_from_2022-03-01_13-37-47.png
Attachment 8: Screenshot_from_2022-03-01_13-37-27.png
Screenshot_from_2022-03-01_13-37-27.png
Attachment 9: Screenshot_from_2022-03-01_13-36-33.png
Screenshot_from_2022-03-01_13-36-33.png
Attachment 10: Screenshot_from_2022-03-01_13-35-11.png
Screenshot_from_2022-03-01_13-35-11.png
Attachment 11: Screenshot_from_2022-03-01_13-34-15.png
Screenshot_from_2022-03-01_13-34-15.png
  346   Tue Mar 1 10:43:36 2022 MS, SHBcheck

'Check Clock Status', 'Check ADC Calibration' and 'Check the White rabbit decoder status': same

 

Attachment 1: Screenshot_from_2022-03-01_11-57-09.png
Screenshot_from_2022-03-01_11-57-09.png
Attachment 2: Screenshot_from_2022-03-01_11-56-49.png
Screenshot_from_2022-03-01_11-56-49.png
Attachment 3: Screenshot_from_2022-03-01_11-56-36.png
Screenshot_from_2022-03-01_11-56-36.png
Attachment 4: Screenshot_from_2022-03-01_11-56-23.png
Screenshot_from_2022-03-01_11-56-23.png
Attachment 5: Screenshot_from_2022-03-01_11-54-23.png
Screenshot_from_2022-03-01_11-54-23.png
Attachment 6: Screenshot_from_2022-03-01_11-53-51.png
Screenshot_from_2022-03-01_11-53-51.png
Attachment 7: Screenshot_from_2022-03-01_11-53-18.png
Screenshot_from_2022-03-01_11-53-18.png
Attachment 8: Screenshot_from_2022-03-01_11-49-32.png
Screenshot_from_2022-03-01_11-49-32.png
Attachment 9: Screenshot_from_2022-03-01_11-49-10.png
Screenshot_from_2022-03-01_11-49-10.png
Attachment 10: Screenshot_from_2022-03-01_11-48-13.png
Screenshot_from_2022-03-01_11-48-13.png
Attachment 11: Screenshot_from_2022-03-01_11-47-31.png
Screenshot_from_2022-03-01_11-47-31.png
  345   Tue Mar 1 10:35:59 2022 MS, SHBnext run on tape server tab

old potentiometer position 19.1 mm

new run number: 83

new potentiometer position : 27.58 mm

 

  344   Tue Mar 1 08:02:26 2022 SHB, MScheck

 

'Check Clock Status', 'Check ADC Calibration' and 'Check the White rabbit decoder status' - no changes

Attachment 1: Screenshot_from_2022-03-01_09-08-41.png
Screenshot_from_2022-03-01_09-08-41.png
Attachment 2: Screenshot_from_2022-03-01_09-08-28.png
Screenshot_from_2022-03-01_09-08-28.png
Attachment 3: Screenshot_from_2022-03-01_09-08-13.png
Screenshot_from_2022-03-01_09-08-13.png
Attachment 4: Screenshot_from_2022-03-01_09-07-55.png
Screenshot_from_2022-03-01_09-07-55.png
Attachment 5: Screenshot_from_2022-03-01_09-07-25.png
Screenshot_from_2022-03-01_09-07-25.png
Attachment 6: Screenshot_from_2022-03-01_09-06-12.png
Screenshot_from_2022-03-01_09-06-12.png
Attachment 7: Screenshot_from_2022-03-01_09-05-55.png
Screenshot_from_2022-03-01_09-05-55.png
Attachment 8: Screenshot_from_2022-03-01_09-05-42.png
Screenshot_from_2022-03-01_09-05-42.png
Attachment 9: Screenshot_from_2022-03-01_09-04-37.png
Screenshot_from_2022-03-01_09-04-37.png
Attachment 10: Screenshot_from_2022-03-01_09-04-04.png
Screenshot_from_2022-03-01_09-04-04.png
Attachment 11: Screenshot_from_2022-03-01_09-03-39.png
Screenshot_from_2022-03-01_09-03-39.png
  343   Tue Mar 1 06:35:42 2022 SHB, MScheck

detector motors are still out

'Check Clock Status', 'Check ADC Calibration' : Clock status test result: Passed 5, Failed 0

'Check the White rabbit decoder status' :  
         Base         Current     Difference
aida01 fault      0xab60 :      0xab8f :      47  
aida02 fault      0x88d3 :      0x8901 :      46  
aida03 fault      0x23aa :      0x23d9 :      47  
aida04 fault      0x15d7 :      0x1606 :      47  
aida05 fault      0x2f03 :      0x2f0f :      12  
White Rabbit error counter test result: Passed 0, Failed 5

Understand the status reports as follows:-
Status bit 3 : White Rabbit decoder detected an error in the received data
Status bit 2 : Firmware registered WR error, no reload of Timestamp
Status bit 0 : White Rabbit decoder reports uncertain of Timestamp information from WR

Attachment 1: Screenshot_from_2022-03-01_07-55-08.png
Screenshot_from_2022-03-01_07-55-08.png
Attachment 2: Screenshot_from_2022-03-01_07-54-46.png
Screenshot_from_2022-03-01_07-54-46.png
Attachment 3: Screenshot_from_2022-03-01_07-53-32.png
Screenshot_from_2022-03-01_07-53-32.png
Attachment 4: Screenshot_from_2022-03-01_07-53-13.png
Screenshot_from_2022-03-01_07-53-13.png
Attachment 5: Screenshot_from_2022-03-01_07-51-22.png
Screenshot_from_2022-03-01_07-51-22.png
Attachment 6: Screenshot_from_2022-03-01_07-47-38.png
Screenshot_from_2022-03-01_07-47-38.png
Attachment 7: Screenshot_from_2022-03-01_07-45-11.png
Screenshot_from_2022-03-01_07-45-11.png
Attachment 8: Screenshot_from_2022-03-01_07-42-33.png
Screenshot_from_2022-03-01_07-42-33.png
Attachment 9: Screenshot_from_2022-03-01_07-41-35.png
Screenshot_from_2022-03-01_07-41-35.png
Attachment 10: Screenshot_from_2022-03-01_07-40-48.png
Screenshot_from_2022-03-01_07-40-48.png
Attachment 11: Screenshot_from_2022-03-01_07-39-56.png
Screenshot_from_2022-03-01_07-39-56.png
  342   Tue Mar 1 06:04:04 2022 SHB, MSpositions

just FYI: current positions saved in files are x=1000 and y=-750. this information is good in case of crash.

 

ELOG V3.1.4-unknown